40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3779     Entry time: Mon Oct 25 23:10:06 2010
Author: Koji 
Type: Update 
Category: IOO 
Subject: MC is now flashing 

[Suresh / Koji]

The MC mirrors are aligned. Now the flashing of the resonances are visible on the MC2 CCD

although the modematching seemed pretty poor.


- The incident power was adjusted to be ~20mW by rotating HWP after the laser source.
The power before the window of the chamber was ~450mW. Where are those missing 1.5W?

- We checked the spot on the last two steering mirrors and the incident beam on MC1.
The beam was too much off from the center of the 1st steering mirror. It was also hitting 1cm north of the MC1.
We adjusted the steering mirrors such that the incident and reflected beams are symmetrically visible at the MC1 tower.

- The MC mirrors are aligned. We first tried to use only MC2 and MC3. And then we used MC1 too as the spot on the MC2 was too high.

- We saw some TEM00 flashes but with many other modes flashing. We checked the beam diameter on the PSL table and on the MC REFL.
The latter one looked twice large as the former one. We concluded the beam is diverging.

- We closed the tank and decided to work on the mode matching tomorrow.

ELOG V3.1.3-